# Full Line Code Completion: Bringing AI to Desktop

### v1

@@ 3.2. Data Pre-processing
Removing comments. Since we are focused on code completion, we filtered out all comments. Moreover, our solution is limited in terms of model size, so for us, it is especially crucial to focus on a single data domain and learn the structure of the source code itself rather than the natural language that describes it. Additionally, we removed empty lines and trailing whitespaces since they do not contribute to source code logic.
@@
Without this step, one will eventually end up with a BPE vocabulary consisting of lots of similar tokens that differ only in the number of tabs in the beginning. The examples of such cases are:
@@
Such an approach allows not only to save space in the BPE vocabulary but also to better handle the token healing problem that is described in Section 3.4.1.
@@
Import dropout. The final step in preprocessing the data for training involves adapting to the workflow in IDEs, in particular, to how developers handle code and imports. Developers often write code first and add imports later, a process facilitated by auto-import features in IntelliJ-based IDEs — they automatically recognize code from external libraries and add the necessary imports. During training, it is crucial to mimic this behavior by teaching the model to predict tokens from libraries even before they are imported. To achieve this, we randomly remove each import statement from training files with a 50% probability, helping the model to learn to anticipate the necessary imports without their explicit inclusion.
@@
We used to have an additional step while prototyping FLCC — formatting all the code (both on training and inference steps) with the Black formatter (Foundation, 2024), which is a popular strict formatter for Python. On the one hand, such procedure allows to simplify stylistic patterns in the code, so it is easier to train the model towards better quality. On the other hand, we have discovered that such strict pre-processing leads to a massive problem where generated suggestions do not match the code style of the user’s project. Moreover, such formatting requires code in the end user’s file to be fully correct, which is not always the case when invoking code completion. Finally, having such a formatting feature requires reimplementing Black in Kotlin so it can be shipped to the end user as a part of FLCC functionality. Thus, we removed this step.
@@ 3.3. Model Training
3.3.1. Tokenization
In FLCC, we use Transformer-based models for the source code, so we chose to use the byte-pair encoding (BPE) (Sennrich et al., 2015) as a standard approach for such models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023).
Character-pair encoding. The traditional BPE is supposed to be used for bytes directly, so it is applicable to all languages if we choose an encoding like UTF-8. However, we decided to use character-pair encoding with slight modifications. Compared to natural language texts, source code primarily consists of common statements or sequences of more than one token (Shin et al., 2019). Some good examples of such recurring “idioms” in Python are
@@ 3.3.1. Tokenization
We chose the vocabulary size of 16,384 tokens because it allows to achieve a decent model quality while maintaining an acceptable memory footprint: for our sizes of models, a significant part of the model’s weights depends on the vocabulary size.
@@
3.3.2. Constructing training examples
@@
We construct each training example as follows. Firstly, we sample a file from the train set. Then, we concatenate file extension, <LANG_SEP_CHAR>, file path, <METAINFO_SEP_CHAR>, and the part of the code above caret, such that the total amount of tokens in the combined three parts fits the maximum context size of the model. Such meta information is always present in the context given to the model, like a header.
3.3.3. Base model
Given their advantages described in Section 2.2.2, we trained a GPT-2-like (Radford et al., 2019) model for FLCC release in PyCharm Pro and DataSpell 2023.3. We experimented with several architectures with self-attention blocks and ended up with a GPT-2-like model since it performed well in terms of the model’s quality, given FLCC restrictions. Additionally, this model can be conveniently converted into ONNX format and has a fast and easy-to-use implementation of the multiplatform inference runtime engine — ONNX Runtime (Microsoft, 2024b). Meanwhile, during the course of 2023, LLaMA architecture and its corresponding llama.cpp (lla, 2024) runtime implementation became popular, so we started its adoption too and shipped LLaMA-based solution to IntelliJ IDEA nightly users in the fall of 2023 and to PyCharm Pro and WebStorm EAP users in the winter of 2024.
@@ 3.3.3. Base model
We use the HuggingFace (Huggingface, 2022) implementations to train GPT-2 and LLaMA models on the described dataset. The models are trained in an automatic mixed precision mode (NVIDIA, 2022) for several days on 8 NVidia A100 GPUs. Following recent advancements in choosing the number of training steps given model and dataset sizes, we used HuggingFace’s approximation (Huggingface, 2024c) for every trained model. For the GPT-2 model, we used the maximum context length of 384 since it performed the best according to our benchmarks, given memory footprint restrictions from target IDEs. Unlike in many similar circumstances, we do not employ knowledge distillation (Sanh et al., 2019), because we discovered that for our sizes of the models, this technique does not provide any quality improvement.
In addition to the standard cross-entropy loss function, we use Recall at 1 (R@1), Recall at 5 (R@5) and Mean Reciprocal Rank at 5 (MRR@5) to control overfitting and perform the primary comparison of models. R@k metrics represent which ratio of the correct suggestions ended up in top k predictions of the model, whereas the MRR@k metric represents the overall quality of top k predictions, taking into account their correctness and position. It is important to remember that these metrics are single-token since language models only learn to predict the next token given context, and the full line is generated by invoking it several times in a row.
3.3.4. Quantization
@@
There are memory footprint and speed constraints among restrictions for FLCC. Thus, to compress the model’s size and speed up its inference, a well-known technique called quantization (Gholami et al., 2022) was used. Quantization refers to changing the way the weights of the model are stored and changing precision that is used for model-related computations. For our GPT-2-like model, we chose ONNX as a format for storing since it allows to easily quantize the model’s weights from FP32 (32-bit floating point numbers) to INT8 (8-bit integer numbers) and to use the ONNX Runtime (ONNX RT) engine. This transformation decreases the size of the model several times, trading it off to several percent of quality loss.
@@
The quantization allowed us to achieve two major goals. Firstly, the size of the model decreased from almost 400 MB to slightly more than 100 MB, which fits within the desired restrictions. Secondly, ONNX RT allows us to shorten the CPU inference time due to more efficient usage of the registers for INT8 data: e.g., the quantized model runs 1.4 times faster compared to full precision model on M2 CPU on MacOS and 2.4 faster on Intel i9 CPU on Linux.
In our ongoing development process, we focus on LLaMA and its corresponding llama.cpp runtime engine because this architecture works faster than GPT-2 with longer contexts, according to our benchmarks. The llama.cpp runtime also allows the trained model to be converted to the required format, quantized, and run on CPU, except in this case, the model’s architecture and the underlying inference engine implementation differ. For LLaMA models, the memory footprint is decreased four-fold too, similar to the ONNX setup for GPT-2. Our recent studies show that quantization of LLaMA allows to achieve 1.4 times higher speed on M2 CPU on Mac and 1.7 times higher speed on Intel i9 CPU on Linux.
@@
In summary, for PyCharm Pro and DataSpell 2023.3 releases, an INT8-quantized GPT-2-like model with 100M parameters was used. The model utilizes a modified BPE tokenizer implemented on top of the YouTokenToMe library. This model scored 0.68 R@1, 0.85 R@5, and 0.75 MRR@5 on the test dataset, with a standard deviation of approximately 0.0001 for each metric. Further Full Line Code Completion releases will use LLaMA-like models because they work better with longer contexts.
3.4. Sequence Generation Algorithm
@@
3.4.1. Token healing
In real-world code completion, the caret position can arbitrarily divide a token, leading to tokenization being inconsistent with the model’s training. An example case is presented in Figure 4. Here, code completion is invoked after “for i”, when the programmer intended to write “for i in range(count):”, and this creates a tokenization issue. During training, the model sees this as a single “for i in range(” token. But in this context, this might be split into “for ” and “i”, differing from the training scenario and reducing the generation quality. This problem escalates with longer average token lengths, increasing the likelihood of mid-token caret positions.
@@ 3.4.1. Token healing
For the example in Figure 4, starting from “for i” and moving left until the start of the line, the prefix matches the vocabulary token “for i in range(”. During code suggestion, the process begins with an empty prefix, discarding hypotheses that do not align with the “for i” prefix. This method, called token healing, resolves generation issues from mid-token positions by reverting to one token prior and constraining the initial generation to the incomplete token.
Figure 4. Token healing. When the caret falls into the middle of a BPE token, we go left until its start (top) and then filter out hypotheses that do not start with the prefix (bottom).
@@ 3.4.2. Collecting hypotheses
Insight 7: Collecting end-of-line-finished hypotheses allows to suggest better completion candidates.
3.4.3. Dynamic number of beam search iterations
@@
Insight 8: A dynamic approach in beam search with carefully selected stopping criteria allows to generate longer suggestions while keeping the mean computation time almost the same.
@@
3.5. Plugin
The Full Line Code Completion plugin for the IntelliJ Platform is the place where the described code completion happens. It implements all the necessary steps: starting from the invocation of code completion and model inference and finishing with the post-processing and showing of the suggestion. The model training pipeline is implemented in Python, so it is not possible to directly reuse code for the plugin, which is written in Kotlin. Thus, we ensure that the processes required for both stages are implemented to be as similar to each other as possible (Zinkevich, 2023). This includes:
  • Data processing. In the plugin, the code from the currently opened file in the editor is taken along with its path and extension. The same Formatter code is applied to it, and the context for the model is composed in the same manner as described in Section 3.2.
  • Sequence generation algorithm. The algorithm described in Section 3.4 is implemented in Kotlin for the plugin.
It is important to remember that the quality of the FLCC feature is predicated not only on the model’s quality but also on various plugin-related features and UX decisions. This section describes our engineering efforts comprising FLCC’s high quality and speed.
3.5.1. Inference
Full Line Code Completion is a computationally-intensive plugin, because it involves the execution of a neural network. To achieve the best possible efficiency and speed, we employed the latest developments of inference engines for Transformer-based models — native CPU runtime for model inference. The implementation takes the following form:
@@ 3.5.1. Inference
  • The server is a wrapper for the model inference runtime: ONNX RT or llama.cpp, depending on the model. All model-related inference computations are managed by the server.
  • gRPC (Google, 2024) protocol is used to establish the connection between the plugin and the server.
In the early fall of 2023, we compared ONNX RT and llama.cpp with our native server implementation on MacOS with an M2 Max processor. The benchmark showed that the mean time from the invocation of completion to the showing of the suggestion with ONNX RT was approximately 75 ms. For llama.cpp, the same procedure took approximately 100 ms. This was the primary reason we focused on an ONNX-based solution for 2023.3 release. Meanwhile, llama.cpp runtime continued to be developed, and the same comparison dramatically changed at the end of November of 2023: it became possible to infer the same LLaMA model in just 50 ms instead of 100 ms, so our development focus changed accordingly.
@@
The last distinguishing feature of our implementation of inference is the caching of the model’s hidden state. The beam search algorithm that we use consists of two steps: context processing and iterative generation. Context processing, which computes token interrelations, is observed to be 3–10 times slower than generation (influenced by hardware and numerical precision).
To optimize the algorithm, we initialize the model at 50% maximum context capacity, enhancing generation speed in two cases:
@@
To match the current context with the cached one, we maintain a list of integers corresponding to cached tokens. This way, we can easily check if the intersection of the contexts is cached or not.
@@
Insight 10: Sophisticated caching strategies allow to speed up code completion functionality without significant quality loss.
3.5.2. Post-processing
@@
Post-processing mostly targets the second goal: after the initial set of hypotheses is generated, we should decide whether there is a useful suggestion in it, and if there is, which one it is.
@@
  • Short suggestions. Excludes single-token entries (already addressed by standard completion) and suggestions shorter than 8 symbols lacking at least two alphanumeric symbols separated with space or punctuation.
  • Safety. Removes dangerous code (e.g., rm -rf, DROP DATABASE), secrets (emails, access keys, etc. (Jackson, 2020)), and profanity.
@@
The Incorrect code filter is the slowest one since it involves running code inspections, and thus it is launched the last and is restricted with maximum execution time. Suggestions with undefined correctness status are not shown to the user. Our research during the preparation for the 2023.3 PyCharm Pro release showed that FLCC benefits from filtering out incorrect code. Only 1% of valuable suggestions are lost, and as a trade-off for such a negligible loss, we establish the distinguishing feature of Full Line Code Completion — suggesting only valid code.
Transformations. After redundant suggestions are filtered out, we are left with a set of “surviving” hypotheses, among which we pick the one with the highest beam search score. The chosen suggestion might still need to be tweaked a little bit, e.g., if not all paired symbols like brackets and quotes are correctly closed. For this reason, as an additional step before showing the code suggestion, we fix such incomplete code closures.
3.6. User Experience
As mentioned before, seamless integration within the IDE is a primary objective for Full Line Code Completion. In pursuit of this UX goal, we have integrated inline completion functionality directly into the IntelliJ Platform API. This allows us to create a proper integration with the standard IntelliJ completion, JetBrains AI Assistant (JetBrains, 2024a), and third-party code completion providers. By January 2024, Salesforce, eBay, Meta, and GitLab have already started using our API, and we continue negotiating it with other providers. Using the same API allows to avoid visual conflicts between providers and to implement chain calls: if one provider cannot suggest code in the given position, another one tries to do so.
@@ 3.6. User Experience
Our plugin also shares some of the features with standard code completion, like auto-import, as described in Section 3.2, and incorrect code detection, as described in Section 3.5.2.
Last but not least, our inline completion implementation supports several hacks and tricks to ensure a smooth user experience. This includes: (1) seamless overtyping, i.e., making sure the suggested code does not blink when you type; (2) computation delay, so the developer does not accept a sudden code suggestion with the TAB keystroke while intending to increase indentation; and others.
Overall, our approach to implementing Full Line Code Completion allows us to cover all restrictions and design principles described at the beginning of this section, meaning that it is local-based, fast, compact, and properly integrated into the IDE.
4. Evaluation
In order to iteratively improve Full Line Code Completion, we established several feedback loops, including offline and online evaluation. Offline evaluation utilizes only a dataset with source code and is frequently used in research because the academic community usually does not have access to a large user base. Online evaluation involves collecting data from real users and is more likely to be used for products. In FLCC, we first check the pipeline quality using offline evaluation, however, we pay a lot more attention to online evaluation. For this reason, in this section, we briefly describe the setup of offline evaluation and give one example of how we used it to evaluate a strategy of collecting hypotheses, and then focus more on the online evaluation.
4.1. Offline Evaluation
When training our model, we calculate recall and mean reciprocal rank, as described in Section 3.3. However, these metrics are single-token and evaluate only the quality of the model, rather than that of the entire code completion pipeline, which is what is actually needed to test the feature.

### v2

@@ III-B Data Pre-processing
Removing comments. Since we are focused on code completion, we filtered out all comments. Because our solution is limited in terms of model size, it is especially crucial to focus on a single data domain and learn the structure of the source code itself rather than the natural language that describes it. Additionally, we removed empty lines and trailing whitespaces since they do not contribute to source code logic.
@@
Without this step, one will end up with a BPE vocabulary consisting of lots of similar tokens that differ only in the number of tabs in the beginning. The examples are:
@@
Such an approach allows not only to save space in the BPE vocabulary but also to better handle the token healing problem that is described further in Section III-D1.
@@
Import dropout. The final step in preprocessing the data for training involves adapting to the workflow in IDEs, in particular, to how developers handle code and imports. Developers often write code first and add imports later, a process facilitated by the auto-import feature in IntelliJ-based IDEs — it automatically recognizes code from external libraries and adds the necessary imports. During training, it is crucial to mimic this behavior by teaching the model to predict tokens from libraries even before they are imported. To achieve this, we randomly remove each import statement from training files with a 50% probability, helping the model to learn to anticipate the necessary imports without their explicit inclusion.
@@
We used to have an additional step while prototyping FLCC — formatting all the code (both on training and inference steps) with the Black formatter [49], which is a popular strict formatter for Python. On the one hand, such procedure allows to simplify stylistic patterns in the code, so it is easier to train the model towards better quality. On the other hand, we have discovered that such strict pre-processing leads to a massive problem where generated suggestions do not match the code style of the user’s project. Moreover, such formatting requires code in the end user’s file to be fully correct, which is not always the case when invoking code completion. Finally, having such a formatting feature requires reimplementing Black in Kotlin so it can be shipped to the end user as a part of FLCC functionality. Thus, we removed this step.
@@ III-C Model Training
III-C1 Tokenization
In FLCC, we use Transformer-based models for the source code, so we chose to use the byte-pair encoding (BPE) [30] as a standard approach for such models [15, 16, 25].
Character-pair encoding. The traditional BPE is supposed to be used for bytes directly, so it is applicable to all languages if we choose an encoding like UTF-8. However, we decided to use character-pair encoding with slight modifications. Compared to natural language texts, source code primarily consists of common statements or sequences of more than one token [50]. Some good examples of such recurring “idioms” in Python are the following:
@@ III-C1 Tokenization
We chose the vocabulary size of 16,384 tokens because it allows for a decent model quality while maintaining an acceptable memory footprint: for our sizes of models, a significant part of the model’s weights depends on the vocabulary size.
@@
III-C2 Constructing training examples
@@
We construct each training example as follows. Firstly, we sample a file from the train set. Then, we concatenate file extension, <LANG_SEP_CHAR>, file path, <METAINFO_SEP_CHAR>, and the part of the code above caret, such that the total amount of tokens in the combined three parts fits the maximum context size of the model. This way, the meta information is always present in the context given to the model, like a header.
III-C3 Base model
Having experimented with several model architectures, including GPT-2 and LLaMA, we ended up with a LLaMA-like model for FLCC for all languages. In our evaluations, LLaMA-like models performed best not only in terms of quality but also speed, thanks to llama.cpp [52] — a multiplatform inference engine for LLaMA-like models.
@@ III-C3 Base model
We use the HuggingFace [53] implementations to train models on the described dataset. The models are trained in an automatic mixed precision mode [54] for several days on 8 NVidia H100 GPUs. Following recent advancements in choosing the number of training steps given the model and dataset sizes, we used HuggingFace’s approximation [55] for every trained model. For our models, we used the maximum context length of 1,536 tokens, since it performed the best according to our benchmarks, given memory footprint restrictions from target IDEs. Unlike in many similar circumstances, we do not employ knowledge distillation [56], because we discovered that for our sizes of the models, this technique does not provide any quality improvement.
To control the model’s overfitting while training, we use a standard cross-entropy loss function evaluated on the validation dataset. However, we pay little attention to well-known metrics like Recall at k (R@k) or Mean Reciprocal Rank at k (MRR@k) because they evaluate only single-token prediction quality, whereas FLCC generates multi-token suggestions.
III-C4 Quantization
@@
There are memory footprint and speed constraints among restrictions for FLCC. Thus, to compress the model’s size and speed up its inference, a well-known technique called quantization [57] was used. Quantization refers to changing the way the model weights are stored and changing the precision that is used for model-related computations. For LLaMA-like models, quantization from FP32 (32-bit floating point numbers) to INT4 (4-bit integer numbers) is supported via llama.cpp. This transformation decreases the size of the model several times, trading it off to a negligible quality loss.
@@
The quantization allowed us to achieve two major goals. Firstly, the size of the model decreased from almost 400 MB to slightly more than 100 MB, which fits within the desired restrictions. Secondly, llama.cpp allows us to shorten the CPU inference time due to more efficient usage of the registers for INT4 data: e.g., the quantized model runs 2 times faster compared to full precision model on M2 CPU on MacOS and 4 times faster on Intel i9 CPU on Linux.
@@
In summary, for Full Line Code Completion, INT4-quantized LLaMA-like models with 100M parameters were used. The models utilize a modified BPE tokenizer implemented on top of the YouTokenToMe library.
III-D Sequence Generation Algorithm
@@
III-D1 Token healing
In real-world code completion, the caret position can arbitrarily divide a token, leading to tokenization being inconsistent with the model’s training. An example case is presented in Figure 4. Here, code completion is invoked after “for i”, when the programmer intended to write “for i in range(count):”, and this creates a tokenization issue. During training, the model sees this as a single “for i in range(” token. However, in this context, this might be split into “for ” and “i”, differing from the training scenario and reducing the generation quality. This problem escalates with larger average token lengths, increasing the likelihood of mid-token caret positions.
@@ III-D1 Token healing
For example, in Figure 4, starting from “for i” and moving left until the start of the line, the prefix matches the vocabulary token “for i in range(”. During code suggestion, the process begins with an empty prefix, discarding hypotheses that do not align with the “for i” prefix. This method, called token healing, resolves generation issues from mid-token positions by reverting to one token prior and constraining the initial generation to the incomplete token.
Figure 4: Token healing. When the caret falls into the middle of a BPE token, we go left until its start (top) and then filter out hypotheses that do not start with the prefix (bottom).
@@ III-D2 Collecting hypotheses
Insight 7: Collecting end-of-line-finished hypotheses allows suggesting better completion candidates.
III-D3 Dynamic number of beam search iterations
@@
Insight 8: A dynamic approach in beam search with carefully selected stopping criteria allows generating longer suggestions while keeping the mean computation time almost the same.
@@
III-E Plugin
The Full Line Code Completion plugin for the IntelliJ Platform is the place where the described code completion happens. It implements all the necessary steps: starting from the invocation of code completion and model inference and finishing with the post-processing and showing of the suggestion. The model training pipeline is implemented in Python, so it is not possible to directly reuse code for the plugin, which is written in Kotlin. Thus, we ensure that the processes required for both stages are implemented to be as similar to each other as possible [58]. This includes:
  • Data processing. In the plugin, the code from the currently opened file in the editor is taken along with its path and extension. The same Formatter code is applied to it, and the context for the model is composed in the same manner as described in Section III-B.
  • Sequence generation algorithm. The algorithm described in Section III-D is implemented in Kotlin for the plugin.
When transferring the prototype from research into the product, it is important to remember that the quality of the feature is predicated not only on the model’s quality but also on various plugin-related features and UX decisions. This section describes our engineering efforts comprising FLCC’s high quality and speed.
III-E1 Inference
Full Line Code Completion is a computationally intensive plugin, because it involves the execution of a neural network. To achieve the best possible efficiency and speed, we employed the latest developments of inference engines for Transformer-based models — native CPU runtime for model inference. The implementation is as follows:
@@ III-E1 Inference
  • The server is a wrapper for the model inference runtime llama.cpp.
  • The server manages all model-related inference computations.
  • gRPC [59] protocol is used to establish the connection between the plugin and the server.
In mid-2023, we compared (1) ONNX Runtime with GPT-2-like models and (2) llama.cpp with LLaMA-like models, both with our native server implementation on MacOS with an M2 Max processor. The benchmark showed that the mean time from the invocation of completion to the showing of the suggestion with ONNX Runtime was approximately 75 ms. For llama.cpp, the same procedure took approximately 100 ms. Meanwhile, llama.cpp continued to be developed, and the same comparison dramatically changed in the late 2023: it became possible to infer the same LLaMA model in just 50 ms instead of 100 ms, so our development focus changed accordingly.
@@
The last unique feature of our implementation of inference is the caching of the model’s hidden state. The beam search algorithm that we use consists of two steps: context processing and iterative generation. Context processing, which computes token interrelations, is observed to be 3–10 times slower than generation (influenced by hardware and numerical precision).
To optimize the algorithm, we initialize the model at 50% maximum context capacity, enhancing the speed of generation in two cases:
@@
To match the current context with the cached one, we maintain a list of integers corresponding to cached tokens.
@@
Insight 10: Sophisticated caching strategies allow to speed up code completion without significant quality loss.
III-E2 Post-processing
@@
Post-processing mostly targets the second goal: after the initial set of hypotheses is generated, we should decide whether it contains a useful suggestion, and if it does, which one it is.
@@
  • Safety. Removes dangerous code (e.g., rm -rf, DROP DATABASE), secrets (emails, access keys, etc. [60]), and profanity.
@@
The Incorrect code filter is the slowest one since it involves running code inspections, and thus it is launched the last and is restricted with maximum execution time. Suggestions with undefined correctness status are not shown to the user. Our research for the 2023.3 PyCharm Pro release showed that FLCC benefits from filtering out incorrect code. Only 1% of valuable suggestions are lost, and as a trade-off for such a negligible loss, we establish the distinguishing feature of Full Line Code Completion — suggesting only valid code.
Transformations. After redundant suggestions are filtered out, we are left with a set of “surviving” hypotheses, among which we pick the one with the highest beam search score. The chosen suggestion might still need to be tweaked a little bit, e.g., if not all paired symbols like brackets and quotes are correctly closed. For this reason, in a final step before showing the code suggestion, we fix such incomplete code closures.
III-F User Experience
As mentioned before, seamless integration within the IDE is a primary objective for Full Line Code Completion. In pursuit of this UX goal, we have integrated inline completion functionality directly into the IntelliJ Platform API. This allows us to create a proper integration with the standard IntelliJ completion, JetBrains AI Assistant [61], and third-party code completion providers. By the fall of 2024, Salesforce, Meta, and GitLab have already started using our API, and we continue negotiating it with other providers. Using the same API allows to avoid visual conflicts between providers and to implement chain calls: if one provider cannot suggest code in the given position, another one tries to do so.
@@ III-F User Experience
Our plugin also shares some of the features with standard code completion, like auto-import, as described in Section III-B, and incorrect code detection, as described in Section III-E2.
Last but not least, our implementation of inline completion supports several hacks and tricks to ensure a smooth user experience. This includes: (1) seamless overtyping, i.e., making sure the suggested code does not blink when you type; (2) computation delay, so the developer does not accept a sudden code suggestion with the TAB keystroke while intending to increase indentation; and others.
Overall, our approach to implementing Full Line Code Completion allows us to cover all restrictions and design principles described at the beginning of this section, meaning that the resulting feature is local-based, fast, compact, and properly integrated into the IDE.
IV Evaluation
In order to iteratively improve Full Line Code Completion, we established several feedback loops, including offline and online evaluation. Offline evaluation utilizes only a dataset with source code and is frequently used in research because the academic community usually does not have access to a large user base. Online evaluation involves collecting data from real users and is more likely to be used for products. In FLCC, we first check the pipeline quality using offline evaluation, however, we pay a lot more attention to online evaluation. For this reason, in this section, we briefly describe the setup of offline evaluation and give one example of how we used it to evaluate a strategy of collecting hypotheses, and then focus more on the online evaluation. It is important to note, however, that we cannot describe all the details of evaluating different models and setups within the scope of this paper, so this section serves to convey the most important results and the overall pipeline of how such features are evaluated in industry.
IV-A Offline Evaluation
When training our model, we calculate recall and mean reciprocal rank, as described in Section III-C. However, these metrics are single-token and evaluate only the quality of the model, rather than that of the entire code completion pipeline, which is what is actually needed to test the feature.